# Mixed birth-death and death-birth updating in structured populations

### v1

@@ 3.3 Fixation and absorption times
    $\mathscr{H}\subseteq\mathscr{G}$ for any sequence $(\mathscr{H}_{\leq N})_{N\geq 1}$ if $\mathscr{H}_{\leq N}\subseteq\mathscr{G}_{\leq N}$ for every $N\geq 1$.
We prove the following.
Theorem 3.4 (Fixation time upper bound).
Let $\delta\in[0,1]$. For all $r^{\star}\geq 1$, there exists $\mathscr{G}^{\star}\subseteq\mathscr{G}$ such that both
  1. 1.
    $\mathsf{FT}^{\delta}_{r}(G_{N})\leq\frac{r}{r-1}\cdot N^{6}$ for all $r\geq r^{\star}$ and all $(G_{N})_{N\geq 1}\in\mathscr{G}^{\star}$, and
  2. 2.
    $\lim\limits_{N\to\infty}\displaystyle\frac{|\mathscr{G}^{\star}_{\leq N}|}{|\mathscr{G}_{\leq N}|}=1$.
In other words, we prove that almost all undirected unweighted graph families have short (polynomial in $N$) fixation time given $r$ is larger than some absolute constant. There is an intuitive characterization of the graph families that reside in $\mathscr{G}^{\star}$: all of the degrees of the vertices are very close to one another. We formally define $\mathscr{G}^{\star}$ in Corollary G.1. For directed graphs, the fixation time can be exponentially long even in the pure Bd case [43, 17].
3.4 Fixation probability estimation algorithm
Next, we turn to the fixation probability for general $r$ and $\delta$. For pure Bd and dB, no known analytical formulas are known in all regimes of $r$. In the absence of an analytical formula, the next best hope is an efficient algorithm that works for all parameter regimes. For Bd, there is a fully polynomial randomized approximation scheme (FPRAS) that outputs $\widehat{\mathsf{fp}}^{\mathrm{Bd}}_{r}(G)$ such that $\mathbb{P}\{|\widehat{\mathsf{fp}}^{\mathrm{Bd}}_{r}(G)-\mathsf{fp}^{\mathrm{Bd}}_{r}(G)|>\varepsilon\cdot\mathsf{fp}^{\mathrm{Bd}}_{r}(G)\}<\alpha$ where $G=(V,E)$ is an undirected unweighted graph on $N$ vertices [1, 43, 46]; this is an algorithm that with confidence $1-\alpha$ outputs an estimate of the fixation probability with multiplicative error $\varepsilon$ in time at most some polynomial of $N$, $1/\varepsilon$, and $\log(1/\alpha)$. Further both Bd and dB (i.e. when $\delta=0,1$), there is a known polynomial time (with respect to $N$) algorithm that computes $\partial\mathsf{fp}_{r}^{\delta}(G)/\partial r\big|_{r=1^{+}}$ for any undirected weighted graph $G$ with $N$ vertices [24]. We give a simple FPRAS for $\mathsf{fp}_{r}^{\delta}(G_{N})$ under the condition that $(G_{N})_{N\geq 1}\in\mathscr{G}^{\star}$. This leads us to the following statement.
Theorem 3.5 (Fixation probability estimation algorithm).
Let $p\in[0,1]$ be a constant. Suppose $r>1$, $\delta\in[0,1]$, $\varepsilon>0$, and $\alpha\in(0,1]$. There exists a polynomial time (with respect to $N$ and $1/\varepsilon$) randomized algorithm $\mathcal{A}$ that given a randomly sampled unweighted undirected Erdős-Rényi graph $G(N,p)$ with probability at least $1-\alpha$ provides an estimate to the fixation probability within a relative error of $\varepsilon$. In other words,
$\mathbb{P}_{\mathbf{G}\sim G(N,p)}\left\{\big|\mathcal{A}(\mathbf{G},r,\delta)-\mathsf{fp}_{r}^{\delta}(\mathbf{G})\big|>\varepsilon\cdot\mathsf{fp}_{r}^{\delta}(\mathbf{G})\right\}<\alpha.$ (1)
The probability is over both the sampling the graph and sampling the randomness for $\mathcal{A}$.
Note that when $\delta=1/2$, we prove (see Theorem 13.1 in the SI) that all undirected unweighted graph families have an FPRAS for $r>1$ using Theorem E.1.
3.5 Fixation probability on cycles
@@
For an undirected cycle $C_{N}$ on $N$ vertices, it is known for Bd that $\mathsf{fp}_{r}^{\mathrm{Bd}}(C_{N})=\mathsf{fp}_{r}^{\mathrm{Bd}}(K_{N})=\frac{1-1/r}{1-1/r^{N}}$ for all $r\in(0,1)\sqcup(1,\infty)$ and $\mathsf{fp}_{r=1}^{\mathrm{Bd}}(C_{N})=1/N$ [30]. For dB, it is known [38, 6, 31] that $\mathsf{fp}_{r=1}^{\mathrm{dB}}(C_{N})=1/N$, and for $r\in(0,1)\sqcup(1,\infty)$ we have,
@@
See Figure 5 for a plot of the fixation probability as a function of $r$ and $\delta$ for a fixed cycle.
3.6 Fixation probability on stars
Figure 6: Fixation probabilities on a star graph with $N=10$ vertices. The various curves represent the various values of $\delta$. The solid lines represent the average of $10^{4}$ simulations of the mixed $\delta$-updating process; the surrounding translucent regions are the $95\%$ confidence intervals. The dashed lines are the theoretical fixation probabilities from Theorem 3.7 (see Appendix J). a, Fixation probabilities as a function of $r$ for $\delta=0,0.25,0.5,0.75,1$ starting with a mutant in the center. The box in the top left corner contains the theoretical fixation probabilities for more $\delta$ values. b, Same as a) except a mutant is initially on a leaf. c, Fixation probabilities as a function of $\delta$ when $r=1$ starting with a mutant in the center (see Theorem 3.7). d, Same as c) except a mutant is initially on a leaf.
Let $S_{N}$ be a star graph with one center labelled $c$ and n:=N1n:=N-1 leaves labelled $\ell_{1},\ldots,\ell_{n}$. For Bd, it is known [4]
@@ 3.6 Fixation probability on stars
For each $i\in\{1,\ldots,n\}$,
Furthermore, we provide an explicit formula for $\mathsf{fp}_{r}^{\delta}(S_{N})$ for all $r>0$.
@@
See Appendix J for the explicit formula for a star for any $r$ and $\delta$. See Figure 6 for a plot of the fixation probability as a function of $r$ and $\delta$ for a fixed star.
3.7 Fixation probability on bidegreed graphs
Figure 7: Examples of $(d_{1},d_{2})$-bidegreed graph families (a-f) and their sensitivities to initial mutant location at $\delta=0,1$ (g-j). Vertices with degree $d_{1}$ are represented as circles whereas vertices with degree $d_{2}\geq d_{1}$ are represented as squares. For the examples depicted, $d_{2}$ is strictly greater than $d_{1}$, but it is also possible for $d_{2}=d_{1}$ per Definition 3.2. Let $n_{1}$ denote the number of vertices with degree $d_{1}$ and $n_{2}$ denote the number of vertices with degree $d_{2}$. The value $\rho_{i}(\delta)$ is the derivative of $\mathsf{fp}_{r=1}(G,\{u\})$ with respect to $\delta$ where $u$ is a vertex with degree $d_{i}$. a, a star graph, $S_{N}$; the $n_{2}=1$ center vertex has degree $d_{2}=6$ while the $n_{1}=6$ periphery vertices have degree $d_{1}=1$ for a total of $N=n_{1}+n_{2}=6+1=7$ vertices. b, a path graph, $P_{N}$; the intermediate vertices have degree $2$ while the endpoint vertices have degree 11. c, a bipartite-biregular graph; the vertices on the left side have degree $\ell=3$ while the vertices on the right side have degree $r=2$; there are $N=10$ vertices. d, a fan graph $F_{N}$; the center vertex has degree $n=2b=6$ while each of the $2$ vertices on the $b$ blades have degree 22, for a total of $N=2b+1=7$ vertices. e, a wheel graph $W_{N}$; the center vertex has degree $n=8$ while the $n$ vertices on the spoke ends have degree 33, for a total of $N=n+1=9$ vertices. f, a book graph $B_{N}$; each of the $p=3$ pages has 22 page corners with degree 22; the 22 seam endpoints have degree $p+1$, for a total of $N=2p+2=8$ vertices. g, A contour plot in the $(\eta,\kappa)$ space where $\eta=n_{1}/n_{2}$ and $\kappa=d_{2}/d_{1}$ are dimensionless parameters. The lines with shapes denote the values of $\eta=n_{1}/n_{2}$ and $\kappa=d_{2}/d_{1}$ for the various bidegreed graphs in a-f. Note that the path graph in our examples since it has $n_{1}\leq n_{2}$ for $N\geq 4$; thus as NN increases, $\eta$ decreases. The intensity of the colorbar denotes the magnitude of $\rho^{\prime}_{1}(\delta=0)$, h, $\rho^{\prime}_{2}(\delta=0)$, i, $\rho^{\prime}_{1}(\delta=1)$, and j, $\rho^{\prime}_{2}(\delta=1)$. The following statements are proven in Appendix K.2: If $d_{1}=d_{2}$, then $\rho_{i}$ is constant with respect to $\delta$. Now suppose $d_{1}<d_{2}$. Then $\rho_{1}$ is strictly decreasing and $\rho_{2}$ is strictly increasing. Further, if $n_{1}>n_{2}$, the function $\rho_{i}$ has its absolute derivative (i.e. sensitivity) maximum at $\delta=1$ and minimum at $\delta=0$. If $n_{1}<n_{2}$, the function $\rho_{i}$ has its sensitivity maximum at $\delta=0$ and minimum at $\delta=1$. If $n_{1}=n_{2}$, then $\rho_{i}^{\prime\prime}(\delta)=0$.
Definition 3.2 (Bidegreed graph).
An unweighted undirected graph $G=(V,E)$ is bidegreed if there exists $d_{1}\leq d_{2}$ such that for all $u\in V$ it is the case that $\deg(u)\in\{d_{1},d_{2}\}$. If $G$ is bidegreed then it must be the case that the pair $(d_{1},d_{2})$ is unique, thus we can also refer to such graphs as $(d_{1},d_{2})$-bidegreed.
Bidegreed graphs are unweighted undirected graphs that have a vertex cut such that the vertices in the cut all have the same degree and the vertices outside the cut have the same degree. Such graphs are encapsulate an alternative notion of close to regular. A $(d_{1},d_{2})$-bidegreed graph is also $(d_{2}/d_{1})$-almost regular. We give a formula for the fixation probability on a bidegreed graph in neutral evolution for any $\delta$.
Theorem 3.8.
Let $G=(V,E)$ be a $(d_{1},d_{2})$-bidegreed graph. Then for all $\delta\in[0,1]$ and $u\in V$,
$\mathsf{fp}_{r=1}^{\delta}(G,\{u\})=\frac{B(\deg(u))}{\sum_{v\in V}B(\deg(v))}$ (9)
where $B\colon\mathbb{N}\to\mathbb{R}$ is defined as
$B(d)=\begin{cases}1&\text{if }d=d_{1},\\ \frac{(1-\delta)d_{1}+\delta d_{2}}{(1-\delta)d_{2}+\delta d_{1}}&\text{if }d=d_{2}.\end{cases}$ (10)
The proof is found in Appendix K.1.

### v2

@@ 3.1 Neutral evolution
    if $n_{1}<n_{2}$, $s_{i}$ has its maximum at $\delta=0$ and minimum at $\delta=1$;
  3. 3.
    if $n_{1}=n_{2}$, then $s_{i}$ is constant with respect to $\delta$.
See Figure 3 for examples of bidegreed graphs and their behaviors with respect to $\delta$. A star graph is a canonical example of a bidegreed graph that has extreme degrees (see Figure 4).
Figure 3: Examples of $(d_{1},d_{2})$-bidegreed graph families (a-f) and their sensitivities in neutral evolution to initial mutant location at $\delta=0,1$ (g-j). Vertices with degree $d_{1}$ are represented as circles whereas vertices with degree $d_{2}\geq d_{1}$ are represented as squares. For the examples depicted, $d_{2}$ is strictly greater than $d_{1}$, but it is also possible for $d_{2}=d_{1}$ per Definition 3.1. Let $n_{1}$ denote the number of vertices with degree $d_{1}$ and $n_{2}$ denote the number of vertices with degree $d_{2}$. The value $\rho_{i}(\delta)$ is the derivative of $\mathsf{fp}_{r=1}^{\delta}(G,\{u\})$ with respect to $\delta$ where $u$ is a vertex with degree $d_{i}$. a, a star graph, $S_{N}$; the $n_{2}=1$ center vertex has degree $d_{2}=6$ while the $n_{1}=6$ periphery vertices have degree $d_{1}=1$ for a total of $N=n_{1}+n_{2}=6+1=7$ vertices. b, a path graph, $P_{N}$; the intermediate vertices have degree 22 while the endpoint vertices have degree $1$. c, a bipartite-biregular graph; the vertices on the left side have degree $\ell=3$ while the vertices on the right side have degree $r=2$; there are $N=10$ vertices. d, a fan graph $F_{N}$; the center vertex has degree $n=2b=6$ while each of the $2$ vertices on the $b$ blades have degree $2$, for a total of $N=2b+1=7$ vertices. e, a wheel graph $W_{N}$; the center vertex has degree $n=8$ while the $n$ vertices on the spoke ends have degree $3$, for a total of $N=n+1=9$ vertices. f, a book graph $B_{N}$; each of the $p=3$ pages has 22 page corners with degree 22; the 22 seam endpoints have degree $p+1$, for a total of $N=2p+2=8$ vertices. g, A contour plot in the $(\eta,\kappa)$ space where $\eta=n_{1}/n_{2}$ and $\kappa=d_{2}/d_{1}$ are dimensionless parameters. The lines with shapes denote the values of $\eta=n_{1}/n_{2}$ and $\kappa=d_{2}/d_{1}$ for the various bidegreed graphs in a-f. Note that the path graph in our examples since it has $n_{1}\leq n_{2}$ for $N\geq 4$; thus as NN increases, $\eta$ decreases. The intensity of the colorbar denotes the magnitude of $\rho^{\prime}_{1}(\delta=0)$, h, $\rho^{\prime}_{2}(\delta=0)$, i, $\rho^{\prime}_{1}(\delta=1)$, and j, $\rho^{\prime}_{2}(\delta=1)$.
Key to our proof of Theorem 3.4 is an explicit formula for the fixation probability on a bidegreed graph for any $\delta$. The proof of the following theorem is in Appendix E.
Theorem 3.5 (Bidegreed fixation probabilities).
On a bidegreed graph with $d_{1}\leq d_{2}$ and for all $\delta$, the fixation probability of a neutral mutant starting on a vertex with degree $d_{1}$ is proportional to $1$. The fixation probability of a neutral mutant starting on a vertex with degree $d_{2}$ is proportional to $\frac{(1-\delta)d_{1}+\delta d_{2}}{(1-\delta)d_{2}+\delta d_{1}}$.
A star graph $S_{N}$ is a canonical example of a bidegreed graph. With center $c$ and n=N1n=N-1 leaves $\ell_{1},\ldots,\ell_{n}$, Theorem 3.5 gives a formula for the fixation probability for any $\delta$. We can also state it explicitly.
Theorem 3.6 (Fixation probability on stars, neutral evolution).
For a star $S_{N}$ with center $c$ and n=N1n=N-1 leaves, for each $i\in\{1,\ldots,n\}$ and all $\delta\in[0,1]$,
The proof is in Section 7 in the SI.
Figure 4: Fixation probabilities on a star graph with $N=10$ vertices in neutral evolution. The various curves represent the various values of $\delta$. The dashed lines are the theoretical fixation probabilities from Theorem 3.6. The translucent regions represent the $95\%$ confidence intervals for the average of $10^{4}$ simulations of the mixed $\delta$-updating process; a, Fixation probabilities as a function of $\delta$ when starting with a mutant in the center. b, Same as a) except a mutant is initially on a leaf.
Similar to fixation probabilities, fixation times in the mixed $\delta$-updating process can exhibit a large range of behaviors relative to $\delta$ (see Figure 5). For graph families $(G_{N})_{N\geq 1}$, it is known $\mathsf{AT}_{r=1}^{\textrm{Bd}}(G_{N})$ is $O(N^{6})$ [1]. For pure dB, it is known $\mathsf{AT}_{r=1}^{\textrm{dB}}(G_{N})$ is $O(N^{5})$ [48]. Relating fixation times to absorption times, the law of total expectation on the absorption time gives the bound $\mathsf{FT}_{r=1}^{\delta}(G,S_{0})\leq\mathsf{AT}_{r=1}^{\delta}(G,S_{0})/\mathsf{fp}_{r=1}^{\delta}(G,S_{0})$ for any $S_{0}\subseteq V$ and $\delta\in[0,1]$.
Figure 5: a, Fixation times, $T$, of various graphs under mixed $\delta$-updating for $N=7$ in neutral evolution. For each graph depicted, the small star symbol on the differently shaded vertex indicates the initial mutant location. i) The fixation time increases and then decreases relative to $\delta$. ii) The fixation time decreases and then increases relative to $\delta$. iii) The fixation time always increases relative to $\delta$. iv) The fixation time always decreases relative to $\delta$.
We prove that the absorption and fixation times are well-behaved in the unbiased process.
Theorem 3.7 (Fixation time).
If $\delta=1/2$ (which is the unbiased case), then the fixation time of a neutral mutant starting in an arbitrary location on any graph of size $N$ is at most $N^{5}/4$.
In our notation, Theorem 3.7 states that for all graphs $G=(V,E)$ on $N$ vertices and any initial mutant location $u\in V$, we have $\mathsf{FT}_{r=1}^{\delta=1/2}(G,\{u\})\leq N^{5}/4$. In the case of $(d_{1},d_{2})$-bidegreed graphs, we are able to say more about their behavior relative to $\delta$. In particular, fixation times on bidegreed graphs are at most some polynomial in $N$, the degree ratio $d_{2}/d_{1}$, and $\delta$.
Theorem 3.8 (Fixation time on bidegreed graphs).
On a bidegreed graph with $d_{1}\leq d_{2}$, the fixation time of a neutral mutant starting on a vertex with degree $d_{1}$ is at most $N^{4}\cdot(d_{2}/d_{1})^{5}/4$. The fixation time of a neutral mutant starting on a vertex with degree $d_{2}$ is at most $N^{4}\cdot(d_{2}/d_{1})^{6}/4$.
For both Theorem 3.7 and Theorem 3.8, we also provide the absorption times starting from an arbitrary location and more precise bounds in the proofs in Appendices F.1 and F.2, respectively.
3.2 Constant selection
Not only do we obtain results for neutral evolution, but also for constant selection. Suppose mutants have relative fitness $r>1$ and wild-types have fitness $1$. We examine two regimes: weak selection ($r>1$ on any graph) and strong selection ($r>(D/d)^{2}$ where $D$ is the maximum degree and $d$ is the minimum degree of the graph). Despite the fact that fixation probabilities and times can be increasing, decreasing, or non-monotonic in $\delta$, we prove that nearly all graphs have short fixation times and provide an efficient algorithm to estimate fixation probabilities.
3.2.1 Weak selection
For undirected unweighted graph families $(G_{N})_{N\geq 1}$, it is known that for the pure Bd process with $r\neq 1$, $\mathsf{AT}_{r}^{\textrm{Bd}}(G_{N})$ is $o(N^{3+\varepsilon})$ for all $\varepsilon>0$ [49, 50]. For $r=1$, it is known $\mathsf{AT}_{r=1}^{\textrm{Bd}}(G_{N})$ is $O(N^{6})$ [1] and $\mathsf{AT}_{r=1}^{\textrm{dB}}(G_{N})$ is $O(N^{5})$ [48]. For any undirected unweighted graph $G=(V,E)$ on $N$ vertices when $r>1$ and $\delta=0,1$, we have $\mathsf{fp}_{r}^{\delta}(G)\geq\mathsf{fp}_{r=1}^{\delta}(G)=\Omega(N^{-2})$; thus $\mathsf{FT}_{r}^{\mathrm{Bd}}(G_{N})$ is $o(N^{5+\varepsilon})$ for all $\varepsilon>0$. Similar to fixation times in neutral evolution, we prove a polynomial bound for any mixed $\delta=1/2$ process.
Theorem 3.9 (Fixation time under weak selection).
If $\delta=1/2$ and $r>1$, then the fixation time of a mutant starting in an arbitrary location on any graph of size $N$ is at most $N^{5}\cdot\frac{r}{r-1}$.
In our notation, Theorem 3.9 states that for any graph $G=(V,E)$ on $N$ vertices and initial mutant location $u\in V$, we have $\mathsf{FT}_{r}^{\delta=1/2}(G,\{u\})\leq N^{5}\cdot r/(r-1)$. The proof is in Section 9 in the SI.
@@ 3.2.1 Weak selection
For an undirected cycle $C_{N}$ on $N$ vertices, it is known for Bd that $\mathsf{fp}_{r}^{\mathrm{Bd}}(C_{N})=\mathsf{fp}_{r}^{\mathrm{Bd}}(K_{N})=\frac{1-1/r}{1-1/r^{N}}$ for all $r\in(0,1)\sqcup(1,\infty)$ and $\mathsf{fp}_{r=1}^{\mathrm{Bd}}(C_{N})=1/N$ [34]. For dB, it is known [42, 6, 35] that $\mathsf{fp}_{r=1}^{\mathrm{dB}}(C_{N})=1/N$, and for $r\in(0,1)\sqcup(1,\infty)$,
@@
See Figure 6 for a plot of the fixation probability as a function of $r$ and $\delta$. See Section 10 in the SI for the proof.
Figure 7: Fixation probabilities on a star graph with $N=10$ vertices when $r\neq 1$. The various curves represent the various values of $\delta$. The solid lines represent the average of $10^{4}$ simulations of the mixed $\delta$-updating process; the surrounding translucent regions are the $95\%$ confidence intervals. The dashed lines are the theoretical fixation probabilities from Theorem 3.11. a, Fixation probabilities as a function of $r$ for $\delta=0,0.25,0.5,0.75,1$ starting with a mutant in the center. b, Same as a) except a mutant is initially on a leaf.
Let $S_{N}$ be a star graph with one center labelled $c$ and n:=N1n:=N-1 leaves labelled $\ell_{1},\ldots,\ell_{n}$. For Bd, it is known [4] that
@@
For a star $S_{N}$ with center $c$ and n=N1n=N-1 leaves, for all $\delta\in[0,1]$ and $r>0$, the fixation probabilities $\mathsf{fp}_{r}^{\delta}(S_{N},\{c\})$ and $\mathsf{fp}_{r}^{\delta}(S_{N},\{\ell_{i}\})$ are given by explicit formulas provided in Section 11 in the SI. Furthermore, when $r=1$ (neutral evolution), these formulas recover Theorem 3.6.
@@
See Figure 7 for a plot of the fixation probability as a function of $r$ and $\delta$ for a fixed star.
3.2.2 Strong selection
For pure Bd and dB, no known closed-form analytical formulas exist for all $r$. For Bd, there is a fully polynomial randomized approximation scheme (FPRAS) that with confidence $1-\alpha$ outputs an estimate of $\mathsf{fp}_{r}^{\mathrm{Bd}}(G)$ with multiplicative error $\varepsilon$ in time polynomial in $N$, $1/\varepsilon$, and $\log(1/\alpha)$ [1, 49, 51]. Furthermore, for both Bd and dB, there is a known polynomial-time (with respect to $N$) algorithm that computes $\partial\mathsf{fp}_{r}^{\delta}(G)/\partial r\big|_{r=1^{+}}$ for any undirected weighted graph $G$ [28]. For graphs with sufficiently large $r$ relative to the degree ratio, we prove the following bounds.
Theorem 3.12 (Fixation probability lower bound under strong selection).
For any graph $G=(V,E)$ on $N$ vertices with maximum degree $D$ and minimum degree $d$, for all $\delta\in[0,1]$ and $r>(D/d)^{2}$, the fixation probability starting at an arbitrary location satisfies $\mathsf{fp}_{r}^{\delta}(G,\{u\})\geq 1/N^{2}$ for all $u\in V$.
Theorem 3.13 (Fixation time upper bound under strong selection).
For any graph $G=(V,E)$ on $N$ vertices with maximum degree $D$ and minimum degree $d$, for all $\delta\in[0,1]$ and $r>(D/d)^{2}$, the fixation time starting at an arbitrary location is at most $N^{7}\cdot\frac{r}{r-1}$.
See Section 12 in the SI for the proofs of the preceding two theorems. Together, these two results show that when $r$ is large enough relative to the degree heterogeneity, the process behaves well. In particular, using the bounds from Theorems 3.12 and 3.13, we give a simple FPRAS for $\mathsf{fp}_{r}^{\delta}(G)$.
Theorem 3.14 (Fixation probability estimation algorithm).
For any graph $G=(V,E)$ on $N$ vertices with maximum degree $D$ and minimum degree $d$, for all $\delta\in[0,1]$, $r>(D/d)^{2}$, $\varepsilon>0$, and $\alpha\in(0,1]$, there exists a polynomial time (with respect to $N$, $1/\varepsilon$, and $\log(1/\alpha)$) randomized algorithm $\mathcal{A}$ that with probability at least $1-\alpha$ outputs an estimate of $\mathsf{fp}_{r}^{\delta}(G,S_{0})$ within multiplicative error $\varepsilon$.
See Section 15 in the SI for the proof. Note that when $\delta=1/2$, the bound on fixation time from Theorem 3.9 gives an FPRAS for all undirected unweighted graphs with $r>1$ (see Theorem 14.1 in the SI).